# Attenuation Bias

### v2

@@ 5 Illustration: Political Knowledge Predicting Duty to Vote
We begin with an example using data from the American National Election Study’s 2024 Time Series Study in order to illustrate our methods, including how our corrected estimator compares to the naive estimator and other approaches. Our latent predictor is political knowledge, which is estimated based on a four-question battery including questions on federal spending, the length of Senate terms, and party control of the House and Senate. [^13] For the full text of items, see Appendix IV.. Many latent traits, such as political knowledge, are often estimated with a small number of indicators in practice. Here, we take the simple approach of estimating the political knowledge level of each respondent with the number of items answered correctly, which we then standardize to have mean 0 and variance 1 within the sample. [^14] To simplify the analysis, we drop all observations with missing values on any of the indicators or the dependent variable, leaving 3,059 of the original 3,349 observations. This standardization makes comparison easier when looking at estimates based on different numbers of indicators as well as different estimation methods, and it also matches the identification restriction used in our derivations above. In applications like this, researchers likely care less about the proportion of correct answers on the specific questions used in the survey and care more about the underlying levels of relative political knowledge held by each respondent, which the question responses provide information about. Given this, identifying the measurement scale through standardization seems a natural approach.
This application demonstrates how our corrected estimators can be applied to any method of estimating a latent trait, including sophisticated models (e.g., item response, factor analytic, etc.) or simpler additive or averaging strategies. Our dependent variable is the response to a 7-point scale question about whether each respondent views voting as a duty or a choice, where we orient the scale so that “Very strongly a duty” is 7 and “Very strongly a choice” is 1. We make no claims here about disentangling the causal relationships between these variables, but only intend this example to illustrate the results of different estimation approaches in commonly encountered situations where latent predictors are estimated based on a small number of indicators.
Figure 3: Grey points show uncorrected linear regression slope coefficients when using various numbers of indicators to estimate the latent trait, then predicting the dependent variable of duty to vote with the point estimates for knowledge. Black dots indicate the averages of these coefficient estimates across all individual estimates using the specified number of indicators to estimate the latent trait. Horizontal lines show the uncorrected OLS estimate $\hat{\beta}_{\hat{X}}$, corrected IV point estimate $\hat{\beta}^{*}_{\text{IV}}$ (which is also equivalent to $\hat{\beta}^{*}_{\text{OLS}}$), and the MOC estimate. See Tables A.IV.3 and A.IV.4 for additional results varying NN in this application.
Naive Estimation (Uncorrected OLS)
We start by fitting a standard linear regression predicting vote duty with estimated knowledge. Here, we are using the point estimates of this latent trait as our predictor without making any corrections to our estimates. In addition to using all four indicators for this estimation, we also separately estimate political knowledge using each possible combination of the four indicators to explore the dynamics of basing our estimates of the latent trait on different numbers of indicators. There are $2^{4}-1=15$ such possibilities since we ignore the case where none of the four questions is included.
Each grey point in Figure 3 plots the point estimate of the slope coefficient from a standard linear regression predicting our dependent variable with estimated knowledge, where knowledge is estimated using a given set of between 1 and 4 of the indicators. The black dots show the average of these OLS slope estimates when using a given number of indicators. We observe that as more items are used to estimate political knowledge, the estimated slope coefficient on political knowledge tends to be higher, which is unsurprising given that a measure using more indicators tends to have less measurement error and, consequently, produce less attenuation than one based on fewer indicators.
The estimate obtained by using all four indicators can be thought of as the naive OLS estimate, since if one were to ignore measurement error, one would presumably use all available indicators to estimate the latent trait, then regress the dependent variable on the point estimates for the latent trait without any further adjustment. This strategy estimates the slope to be 0.42, as indicated in the figure by the horizontal dashed line.
Corrected Split‑Indicator Estimator
We can also create two separate measures of knowledge by partitioning the four indicators into two sets, each including two of the knowledge items. For example, we could define $\hat{X}_{1}$ as the standardized number of correct answers that a respondent obtains on the first two knowledge questions and define $\hat{X}_{2}$ as the standardized number of correct answers that a respondent obtains on the third and fourth knowledge questions. The sample correlation between these two estimates in this case would be 0.25, suggesting a relatively large amount of measurement error. We could then use this correlation to correct the slope coefficient estimate.
The horizontal dotted line in Figure 3 shows the value of our corrected estimator $\hat{\beta}_{\text{IV}}^{*}$, which takes the median of the within-partition averages of the corrected IV estimators over all possible partitions of the four items into two sets of two items. As shown above, this estimator is equivalent to the corrected OLS estimator $\hat{\beta}^{*}_{\text{OLS}}$. This corrected estimate is 0.62, which is nearly 50% larger than the baseline (uncorrected) OLS estimate, showing how accounting for measurement error can yield substantively meaningful changes in our estimates. Note also that weak instruments do not appear to be a problem, given the high correlation between the two sets of latent trait estimates. The $F$-statistic from the first stage averaged well over 200 across all possible partitions, with the smallest $F$-statistic being 134.6—far above the commonly cited threshold of 10 [43].
@@
Had we used the standard (uncorrected) IV estimator (not shown in the figure) we would have estimated the slope as 1.27 on average across these partitions of the indicators, which is more than double the value of corrected estimate. [^15] If we average the uncorrected IV estimates within each partition and then take the median over these within-partition averages, following the strategy for our corrected estimator, the result is 1.14, which is still 85% higher than our corrected estimate. This illustrates the high degree of amplification bias—estimating the slope as being larger in magnitude on average than the true value—that results from applying standard IV-based corrections for measurement error to the latent predictor setting.
@@
We could also have used an item response model to estimate political knowledge in this application instead of the simple additive strategy. This would entail obtaining point estimates (e.g., maximum likelihood parameter values or posterior means) of the political knowledge levels for each respondent, then using those as predictors in either baseline OLS regressions or IV regressions, as above, to which we could apply appropriate corrections.
Using an item response model to estimate our latent traits here produces baseline OLS slope estimate of 0.45 (quite similar to the 0.42 using the standardized proportion correct over the set of 4 items). [^16] Here we use the ideal function in the pscl R package to estimate latent traits. Using our corrected estimator with the item response model estimates, we estimate the slope as 0.67, which is slightly higher than the corrected estimate using an additive measure rather than the item response model. Estimating an item response model rather than using simple aggregation methods, such as the total number correct, may be expected to provide a more precise estimate of the latent trait, which could account for the slightly larger estimates obtained for both the standard OLS and corrected IV approaches.
When using the latent traits model, we can also apply MOC, which is the primary solution proposed in the political science latent traits modeling literature for accounting for the measurement error inherent in latent predictors. Note that MOC requires some sort of model for estimating the latent trait since we need to be able to obtain draws from the posterior distribution for the latent traits or something similar (e.g., some likelihood-based distribution), so it cannot be applied to simple latent trait estimation strategies like adding or averaging that do not have an explicit model.
But as discussed above, MOC averages over a set of draws from the posterior distribution for the latent traits given the indicators, with each of these draws being an even noisier measure of the true latent trait than the baseline strategy of using the point estimates (e.g., posterior means or maximum likelihood point estimates) for the latent traits as our predictor. In our political knowledge example, MOC yields a point estimate of 0.20, as shown by the solid horizontal line in Figure 3. This is less than half the size of the unadjusted OLS approach, which is itself known to be attenuated, and is less than one-third the size of our corrected estimate. This is consistent with most previous applications of MOC in the literature, which tend to produce coefficient estimates that are smaller in magnitude than the (presumably already attenuated) naive estimates treating latent trait estimates as known, and illustrates how, rather than helping, MOC actually exacerbates the central problem of attenuation bias that arises when using latent predictors in regression models.
This example also suggests that the specific method used to estimate a latent trait (e.g., simple aggregation or averaging versus item response models) may often matter much less for the resulting slope coefficient estimates than does the method of adjusting (or not adjusting) for the inherent measurement error in the latent trait estimates. While there are only minor differences when switching from the simpler aggregation approach to using an item response model here, estimates differed by nearly 50% when moving from naive OLS to our corrected estimator, and the differences were even more dramatic for MOC and unadjusted IV.
6 Simulation: The Role of Sample Size & Dimensionality
In this section, we study how slope estimates for a latent predictor vary with the number of indicators ($M$) and the sample size ($N$), and how different estimation strategies compare. We evaluate (i) naive OLS that treats point estimates of $X$ as known, (ii) uncorrected IV, (iii) the method of composition (MOC), and (iv) our correlation‑adjusted estimator (IV/OLS‑equivalent).
We generate binary indicators using the standard probit ideal‑point model
@@ 6 Simulation: The Role of Sample Size & Dimensionality
and relate the (unobserved) trait to an outcome via
$Y_{i}=\beta_{0}+X_{i}\beta_{X}+\epsilon_{i}.$ (7)
As above, $X$ is latent with no objective scale; we identify it by imposing $\mathbb{E}[X]=0$, $V[X]=1$, and a directional restriction.
To obtain parameter values for our simulations, we first analyze data from the 2016 CCES (N=38,851N{=}38{,}851), which includes 30 agree/disagree policy items that we use as indicators to estimate respondent ideology ($X$). Our dependent variable ($Y$) is presidential vote (0=Democratic, 1=Republican; other responses dropped). The estimates from these analyses are used in these simulations (see Appendix III for more detailed information about this setup and the simulation results). In the largest $N$ and $M$ setting, the naive slope is 0.38, our corrected estimator is 0.40, and the ground‑truth value is 0.40. The small naive–corrected gap in this calibration reflects that M=30M{=}30 yields relatively precise measurement; our simulations below systematically vary $M$ and $N$ to show when corrections matter most.
We fix item parameters $(\eta,\alpha)$ and regression parameters $(\beta_{0},\beta_{X},\sigma^{2}_{\epsilon})$ to their CCES point estimates and vary
$N\in\{10^{2},10^{3},10^{4}\},\qquad M\in\{6,12,30\}.$
For each $(N,M)$, we: (1) sample $N$ respondents and $M$ items without replacement to form $\mathbf{W}\in\{0,1\}^{N\times M}$; (2) estimate item parameters and latent $X$ via EM (binIRT in emIRT) [^17] We use binIRT in emIRT [24] with default priors.; (3) form split‑indicator scores and compute the correlation‑based correction; and (4) estimate $\beta_{X}$ using naive OLS, our corrected estimator $\hat{\beta}^{*}_{\text{IV}}$ (equivalently $\hat{\beta}^{*}_{\text{OLS}}$), uncorrected IV, and MOC. We average results over 16 random partitions of items in the two‑stage procedure.
This setup lets us compare bias and variability across methods as information about $X$ changes along two key axes: more indicators (larger $M$) and more observations (larger $N$). It also highlights when our corrected estimator yields the largest—and substantively meaningful—revisions relative to naive OLS, uncorrected IV, and MOC.
@@
Figure 4 presents the results, focusing on how the absolute bias, standard error of estimation, and RMSE of each of these three estimators vary as the sample size and number of policies change. The standard (uncorrected) OLS estimator shows relatively large absolute bias (note that the actual bias is toward zero, but we use absolute bias to facilitate easier comparison of the magnitude of bias between estimators that are biased in different directions). Recall that the true value for the slope coefficient is 0.40, meaning that the absolute bias around 0.10 as seen for $M=6$ represents a 23.81% change from the true parameter value. This bias does not change meaningfully with the number of observations, but does decrease as the number of indicators used ($M$) grows. Our corrected estimator exhibits very low bias across all values of $M$ and $N$. The standard deviation does not differ meaningfully between the uncorrected OLS and corrected estimates, suggesting that the reduced bias of our proposed estimator is the main difference.
In all cases here, MOC produces worse estimates—notably higher bias and RMSE—than not only our corrected estimator but also naive OLS, which ignores measurement error altogether. This suggests that MOC is not a useful method for correcting for measurement error and, in fact, makes things much worse. Finally, the uncorrected IV estimator has the highest absolute bias across all $M$ and $N$ (as discussed above, the bias of uncorrected IV is away from zero, rather than toward zero as with the naive estimator and MOC). Differences between the standard deviation of the three estimators are relatively small, particularly between the uncorrected OLS and our corrected estimator (the uncorrected IV is much more variable than the other two when $M=6$).
Overall, the RMSE indicates that our corrected estimator outperforms both the other estimators considered across a wide range of $M$ and $N$. Although differences in the RMSE decrease with $M$ (and very slightly with $N$), we see that even with 30 indicators, a value that is higher than the number used in many applications, there are still meaningful relative improvements (e.g., with adjustment decreasing RMSE from 0.025 to <<0.01).
Figure 4: Simulation results varying NN and $M$.
@@
Presenting these simulation results in a different way, Appendix Table A.IV.1 shows the percentage change in coefficient estimates when moving from naive OLS to our corrected estimator, for different values of both $N$ and $M$. While these differences are relatively small when using 30 indicators, we see changes around 20% when $M=12$, 25% when $M=10$, and 35% when $M=6$. Note that estimating latent traits from smaller indicator sets is far more common in practice than using 30 or more indicators. This suggests that accounting for measurement error in latent predictors estimated from binary indicators is likely to make a substantial difference in many applications.
In the Appendix (Figure A.III.3), we also perform various comparisons with other methods not shown in Figure 4, such as the multiple over-imputation and Bayesian joint factor–outcome models requiring simultaneous specification of both measurement and structural components and often depend on intensive MCMC sampling [29, 6]. We find that, in general, overimputation performs well with a large number of factors, but struggles somewhat with smaller $M$, while joint modeling performs best among the group, although it never seems to systematically outperform the adjusted approach outlined here. This suggests that simple partitioning of indicators and attenuation adjustment can be an attractive approach, especially as it does not require complex joint posterior inference, and adjustment factors can be reused for new outcomes.
@@
For a power analysis, see Section A.III.2 of Appendix III. There, we find that the corrected estimator provides improved coverage compared to unadjusted approaches. This improvement does not bring coverage to nominal levels for small values of $M$ (where we undercover, possibly due to small finite-sample bias); for large values of $N$ and $M$, we find nominal coverage for large values of $N$ and $M$ for the corrected estimator, but not others.
7 Further Applications: Democracy and Text Sentiment as Latent Predictors
Up to this point, we have provided some theoretical background, an illustrative example of our proposed method, and generated further evidence through simulation. In the political knowledge application, our proposed correction resulted in very large changes to the estimated slope, suggesting that naive linear regression ignoring measurement error produced a high degree of attenuation bias. In the ideology example that motivated our simulations in Section 6, the corrected estimator produced much smaller changes when the latent trait was estimated using all 30 indicators, but when using fewer indicators to estimate the latent trait, as is more common in the literature, our proposed correction produced substantively large changes.
How significant are these measurement error dynamics in practice when attempting to answer different substantive social science research questions? Here we turn to two more substantive applications—one drawn from comparative politics, another from the realm of text analysis—illustrating how our method can be used to assess whether measurement error in a given application appears to bias estimates to a meaningful degree. Both of these examples also use measurement models for the latent traits that do not have full underlying statistical models. This means that while our corrected estimation approach can be used, other methods, such as MOC or full joint estimation, would not work with these measurement strategies.
7.1 Principal Component Analysis: Democracy and Growth
We here re-analyze the relationship between democracy and growth using this measurement error perspective. Specifically, we investigate whether adjusting for the attenuation bias inherent in latent democracy measures materially alters the estimated impact of democratic institutions on subsequent economic growth. To generate measures of democracy, we use the raw data from the Freedom House Index [22]. The raw indicator-level data collected by index creators, available between 2013 and 2022, provide information on a total of 17 sub-indicators related to electoral processes, political pluralism, the functioning of government, freedom of expression, associational rights, the rule of law, and personal autonomy. Here, we employ a simple measurement strategy, estimating the latent variable of democracy with the first principal component of these indicators. This also provides an illustration of how our approach can be used with measurement strategies for the latent trait of interest that do not have an underlying full statistical model, such as principal components analysis. For the dependent variable, we consider the long difference in log GDP per capita between 2022 and 2013.181818The long difference is: $\Delta\log\bigl(\textrm{GDP per capita}\bigr)\;=\;\log\bigl(\textrm{GDP per capita}_{2022}\bigr)\;-\;\log\bigl(\textrm{GDP per capita}_{2013}\bigr),$ which approximates the cumulative growth rate over the period.
Table 1 shows the naive OLS estimate as well as our corrected estimate of the slope coefficient on democracy. These estimates are nearly identical, with our correction resulting in a half-percentage-point change. In this example, then, our methods suggest that measurement error may be a negligible issue. This is because of the high degree of correlation between the indicators used to measure democracy, which suggests that the latent variable is estimated with considerable precision under the assumptions used here. Appendix Section IV shows that even if we use a smaller subset of the 17 indicators to estimate democracy, the impact of measurement error seems relatively small.
Table 1: Comparison of democracy coefficients (outcome: growth), adjusted and unadjusted. Quantile‐based confidence intervals constructed via bootstrap.
Value Estimate (with CI)
Unadjusted Coefficient 0.0255 [-0.0060, 0.0612]
Adjusted Coefficient 0.0257 [-0.0060, 0.0616]
Change in Estimate 0.520%
7.2 Text Analysis: Sentiment and Incumbent Vote Share
We next analyze an application of text analysis in the realm of political communication. Our primary objective is to determine how sentiment in political rhetoric predicts incumbent success in the next election. Following [15], we might expect a positive coefficient on net positivity in regression analysis; however, negative rhetoric has become increasingly common throughout the 2010s, particularly as social media amplified partisan messaging strategies [25]. We focus the analysis on this period.
For data, we use house‐level general–election returns from the MIT Elections Lab. Congressional speech texts are drawn from the Congressional Record; House member partisanship is drawn from Voteview [33]. After identifying incumbents and computing their district-level two-party vote shares, we aggregate each representative’s House floor speeches by speaker and congressional session. Due to OCR-induced inconsistencies in speaker names preventing full matching, we successfully link on average 350 incumbents’ records per year.
We score each House member’s floor speeches with a pre‑trained RoBERTa sentiment model [34] that outputs probabilities over five labels (‘strongly negative” to ‘strongly positive”). Because many speeches exceed the 512‑token limit, we segment texts, compute segment‑level probabilities, map labels to numeric scores, and average across segments to obtain a speaker‑level sentiment measure. Split samples are formed by repeating this procedure on two non-overlapping halves of the corpora. As this measurement does not rely on a formal statistical model, our corrected estimator is straightforward to implement, whereas MOC, MO, and full joint estimation would require bespoke modifications to large‑scale training pipelines.
Table 2: Comparison of sentiment coefficients, adjusted and unadjusted. Quantile-based confidence intervals constructed via bootstrap.
Value Estimate
Unadjusted Coefficient -2.83 [-4.28, -1.31]
Adjusted Coefficient -3.23 [-4.92, -1.48]
Change in Estimate -14.40%
The naive OLS slope coefficient estimate ignoring measurement error is -2.83. as shown in Table 2. After applying the split-sample correction to adjust for attenuation bias, our corrected estimator is -3.23, which represents a -14.40% increase in magnitude. Here, appropriately accounting for measurement error in the latent predictor yields a small but non-trivial change in our estimates.

### v3

@@ 4 Illustration: Political Knowledge Predicting Duty to Vote
To make the identification and measurement-error issues concrete, we use the 2024 ANES to compare four common approaches in a familiar latent-predictor setting: unadjusted OLS, MOC, uncorrected IV, and our correlation-corrected estimator. The latent predictor is political knowledge, measured with a four-item battery (federal spending, Senate term length, and party control of the House and Senate). [^11] For full item wording, see Appendix IV. As is typical for many latent traits applications, the indicator set is small. We estimate each respondent’s knowledge as the number of correct answers (0–4) and then standardize within-sample to mean 0 and variance 1.121212We drop observations with missing values on any indicator or on the dependent variable, leaving 3,059 of 3,349 respondents. This normalization matches the identifying restriction used above and facilitates comparisons across different indicator subsets and estimation methods. Substantively, we interpret the scale as relative knowledge in the sample, rather than the proportion correct on these specific items.
This application also illustrates the modularity of our correction: it can be applied to any latent-trait estimator, from simple additive scores to item-response or factor models. The outcome $Y$ is a 7-point item on whether voting is a duty or a choice, coded so that “Very strongly a duty” is 7 and “Very strongly a choice” is 1. We make no causal claims; the goal is purely to contrast estimation strategies when a latent predictor is measured with few indicators.
Figure 1: Grey points are uncorrected OLS slope coefficients from regressing duty-to-vote on estimated knowledge, where knowledge is constructed from each nonempty subset of the four indicators (1–4 items). Black dots are the mean slope within each indicator-count. Horizontal lines show: the unadjusted OLS estimate using all four items ($\hat{\beta}_{\hat{X}}$), the correlation-corrected estimate ($\hat{\beta}^{*}$), and the MOC estimate (using an item-response model). See Tables A.IV.3 and A.IV.4 for additional results varying NN.
Baseline: Uncorrected OLS
We begin with the standard practice: regress vote duty on the point estimates of knowledge, without correcting for measurement error. In addition to the all-four-items score, we recompute knowledge for every nonempty subset of the four indicators to show how the slope changes as measurement becomes more (or less) precise. There are $2^{4}-1=15$ such subsets.
Figure 1 shows a clear pattern: as more items are used to construct knowledge, the estimated slope increases. This is consistent with attenuation from measurement error: adding indicators reduces noise and increases the slope toward its latent-scale value. Using all four items (the typical baseline choice) yields an unadjusted OLS slope of 0.42 (dashed line).
Correlation-Corrected Estimator
Next, we form two independent estimates of knowledge by splitting the four items into two pairs, estimating $\hat{X}_{1}$ and $\hat{X}_{2}$ separately, and using their correlation to undo identification-aware attenuation. For example, splitting the first two items versus the last two yields $\operatorname{Cor}(\hat{X}_{1},\hat{X}_{2})=0.25$, indicating substantial measurement error. We compute the correlation-corrected slope within each possible 2–2 split and summarize across splits by taking the median of the within-split averages.
The dotted line in Figure 1 shows the resulting estimate, $\hat{\beta}^{*}=0.62$, nearly 50% larger than the unadjusted OLS estimate. Under the IV interpretation of the split-half procedure, weak instruments do not appear to be a concern: the first-stage $F$-statistic averages well above 200 across partitions, with a minimum of 134.6—far above the conventional threshold of 10 [undefak] (high $F$ here is driven by the large sample size, even with only moderate split-half correlation).
@@
For comparison, applying standard (uncorrected) IV to these split measures (not plotted) produces severe amplification: the average slope across partitions is 1.27, more than double our corrected estimate. [^13] If we instead average uncorrected IV estimates within each partition and then take the median across partitions (mirroring our aggregation strategy), the result is 1.14, still 85% larger than our corrected estimate. This illustrates how conventional IV-style measurement-error corrections can over-correct in the latent-predictor setting.
@@
Nothing in our correction depends on using an additive score. Alternatively, we can estimate knowledge with an item response model, obtain respondent-level point estimates (e.g., MLEs or posterior means), and then repeat the same baseline and corrected regressions. [^14] For estimation, we use the ideal function from the pscl package in R. With IRT-based scores, the baseline OLS slope is 0.45 (close to 0.42 under the additive score), while the correlation-corrected estimate is 0.67 (slightly above 0.62). These modest shifts are consistent with IRT producing somewhat more precise trait estimates, but the main differences still come from whether measurement error is corrected at all.
With an explicit latent-trait model in hand, we can also apply MOC, the standard recommendation in much of the applied latent-traits literature for propagating uncertainty into second-stage regressions. A practical limitation is that MOC requires posterior draws (or approximations), so it is not available for ad hoc scores like simple sums or averages. More importantly for our purposes, MOC averages over regressions fit to noisier draws of the latent trait than the point estimates, which can exacerbate attenuation. In this application, MOC yields a slope of 0.20 (solid line in Figure 1), less than half the unadjusted OLS estimate and less than one-third of our correlation-corrected estimate. This pattern mirrors many published MOC applications and underscores that, in latent-predictor regressions, MOC can worsen the central first-order problem: attenuation bias from measurement error.
Overall, this example suggests that at least here the choice of latent-trait estimator (simple aggregation versus IRT) matters less for the slope than how measurement error is handled. Moving from unadjusted OLS to our correction shifts the estimated relationship by roughly 50%, while MOC and unadjusted IV push estimates much farther in opposite directions.
5 Comparing the Performance of Estimators with Simulation
To compare the performance of our correlation-corrected estimator with other approaches, in this section we conduct Monte Carlo simulations, generating simulated data ($W$ and $Y$) with parameter values set to point estimates from real data. Specifically, we rely on the 2016 Cooperative Congressional Election Study (CCES) (NN=38,851), a survey that includes 30 agree/disagree policy items that we use as indicators to estimate respondent ideology ($X$). We assume a commonly used item response model
@@ 5 Comparing the Performance of Estimators with Simulation
and our dependent variable ($Y$) in the regression model is reported presidential vote (0=Democratic, 1=Republican; other responses dropped).
Although the type of data and the approach to estimating the latent trait here are quite common in the literature, both the number of observations ($N$) and a number of indicators ($M$) in the CCES are much larger than in most latent traits applications in political science and related literatures. This allows us to assess the performance of estimators across a range of $N$ and $M$ values, from small to large. Specifically, we set parameter values equal to estimates from the full CCES data. We then simulate data for $N\in\{10^{2},10^{3},10^{4}\}$ and $M\in\{6,12,30\}$, sampling observations and indicators randomly for each value of $N$ and $M$. We then estimate $\beta$ with these simulated datasets using each of several approaches discussed above (see Appendix III for more detailed information about this setup and the simulation results).
@@
Figure 2 presents the results, focusing on how absolute bias, standard error, and RMSE vary with sample size $N$ and number of indicators $M$ for each estimator in the comparison (Uncorrected OLS, Cor. Corrected, Method of Composition, and Bayesian, Full). Generally, the standard deviation does not differ much between the estimators, so differences in RMSE are driven primarily by their relative biases. The standard OLS estimator shows relatively large absolute bias (note that the actual bias is toward zero, but we use absolute bias to facilitate easier comparison of the magnitude of bias between estimators that are biased in different directions). The true value of $\beta$ is 0.40, meaning that the absolute bias around 0.10 as seen for $M=6$ represents a 23.81% change relative to the true parameter value. This bias does not change meaningfully with the number of observations, but does decrease as the number of indicators used grows. Our correlation-corrected estimator exhibits very low bias across all values of $M$ and $N$. In fact, the performance of our estimates is very close to that of the full joint estimation strategy, which is often much more cumbersome (or, when the measurement strategy for the latent trait does not specify a full statistical model, may be infeasible). In all cases here, MOC produces worse estimates—notably higher bias and RMSE—than our correlation-corrected estimator. MOC performs worse than standard OLS, which ignores measurement error altogether.
Figure 2: Simulation results varying NN and $M$. Results show comparison of performance of estimators from simluations described in Section 5.
@@
Presenting these simulation results in a different way, Appendix Table A.IV.1 shows the percentage change in coefficient estimates when moving from standard OLS to our corrected estimator, for different values of both $N$ and $M$. While these differences are relatively small when using 30 indicators, we see changes around 20% when $M=12$, 25% when $M=10$, and 35% when $M=6$. Because estimating latent traits from smaller indicator sets is far more common in practice than using 30 or more indicators, these results for $M$ between 6 and 12 are likely to be the empirically relevant regime. This suggests that accounting for measurement error in latent predictors estimated from binary indicators is likely to make a substantial difference in many applications.
In the Appendix (Figure A.III.3), we also perform various comparisons with other methods not shown in Figure 2, such as the multiple over-imputation and standard (uncorrected) IV. We find that, in general, overimputation performs well with a large number of factors, but struggles somewhat with smaller $M$, while uncorrected IV has by far the highest bias (in this case away from zero) and RMSE of any method considered.
@@
For a power analysis, see Section A.III.2 of Appendix III. There, we find that the corrected estimator provides improved coverage compared to unadjusted approaches. This improvement does not bring coverage to nominal levels for small values of $M$ (where we undercover, possibly due to small finite-sample bias); for large values of $N$ and $M$, we find nominal coverage for large values of $N$ and $M$ for our corrected estimator, but not others.